OpenStudy (anonymous):

The midpoint of A (-4, 2) and B(8, 5) is (2, 7/2) (-6, 3/2) (4, 7/2)

OpenStudy (anonymous):

@austinL @amistre64

OpenStudy (austinl):

Ok, so the midpoint formula is as follows.

OpenStudy (austinl):

\[(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\]

OpenStudy (anonymous):

[6,3/2]

OpenStudy (anonymous):

M(x,y) x=(-4+8)/2=2 y=(2+5)/2=7/2

OpenStudy (anonymous):

Thanks everyone

OpenStudy (austinl):

\[(\frac{-4+8}{2},\frac{2+5}{2})=(\frac{4}{2},\frac{7}{2})\]

OpenStudy (anonymous):

Thanks